# SAMM: Sharded Automated Market Makers

### v1

@@ 4.3 CPMM Trading fee
5 SAMM: Sharded AMM
We introduce SAMM, the first sharded Automated Market Maker. We explain the driving rationale and derive two desired properties of the contract (§5.1). We present the SAMM trading fee function (§5.2) and find parameters to fullfill both properties (§5.3).
5.1 SAMM Structure and Properties
@@
{restatable}
[$c$-Non-Spliting]proppropertyconcavity Let $m\geq 2$. Given a set of output amount by $\{O^{A}_{j}|1\leq j\leq m,O^{A}_{j}>0\}$, denote by $\tilde{O}^{A}$ the sum of the amounts in the set, $\tilde{O}^{A}=\sum_{j=1}^{m}O^{A}_{j}$. For the constant $10<c<$ 1 and the deposited amount of tokens $R^{A}$ and $R^{B}$, if $\frac{\tilde{O}^{A}}{R^{A}}\leq c$ , then the cost of trading $\tilde{O}^{A}$ token A is less than the sum of the cost of trading $O^{A}_{j}$ token A for $1\leq j\leq m$, i.e.,
@@
5.1.2 c-smaller-better property
Our second goal is to maintain balanced volumes in all shards. This is crucial because the volume directly affects the slippage. When there are stark differences in pool sizes, with some being much smaller than others, the slippage in trading within these smaller pools is significantly greater than in larger ones. This discrepancy can result in transactions clustering in the larger pools rather than spread evenly, reducing parallelism. We address this by incentivizing liquidity providers to allocate their tokens to the shards with lower volumes. Intuitively, this ensures that smaller pools receive more frequent fees from traders when the volumes of pools are not balanced, which incentivizes liquidity providers to deposit tokens in these smaller pools. Therefore, the system would converge to the state where all pools have balanced volumes, and traders then randomly select pools for trading. Similar to the $c$-Non-Splitting property, large transactions suffer from high slippage, which leads to a strong advantage of larger pools. So here too, we refine this requirement to scenarios where the ratio of the traded amount and the deposited amount is below a threshold $c$. We call this the $c$-smaller-better property.
{restatable}
[$c$-smaller-better]proppropertysmallerpoolsmallercost Given an output amount ${O^{A}>0}$, for any two pools with deposited token amounts $(R^{A}_{i},R^{B}_{i})$ and $(R^{A}_{j},R^{B}_{j})$, respectively, and $R^{A}_{i}<R^{A}_{j}$. For the constant $10<c<$ 1, if $\frac{O^{A}}{R^{A}_{j}}<\frac{O^{A}}{R^{A}_{i}}\leq c$ and $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}$, then the cost of trading $O^{A}$ token A in the smaller pool is less than that in the larger pool, i.e.,
@@ 5.1.3 $c$ value
Finally, we observe that the traditional CPMM cost function, $G_{\textit{CPMM}}(R^{A},R^{B},O^{A})$, does not satisfy either property (Appendix C).
5.2 Trading Fee Design
To satisfy properties 5.1.1 and 5.1.2, we first generalize the trading fee function to provide flexibility in the incentive design (§5.2.1). Then, we propose a specific trading fee function (§5.2.2).
5.2.1 Generic Trading Fee Function
@@
5.2.2 Bounded-Ratio Trading Fee Function
@@
$\textit{tf}(R^{A},R^{B},O^{A};\beta_{1},\beta_{2},\beta_{3},\beta_{4})\mathrel{\mathop{:}}=\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}\,\,.$
@@
$\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A};\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5})\mathrel{\mathop{:}}=\\ \frac{R^{B}}{R^{A}}O^{A}\times\max\{r_{\min},\min\{r_{\max},\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\}\}$ (10)
The ratio $\frac{R^{A}}{R^{B}}$ represents the market price of token A relative to token B (see Section 4.2). The product $\frac{R^{B}}{R^{A}}O^{A}$ is thus the trader’s net payment in terms of token B without slippage. Then $\max\{r_{\min},\min\{r_{\max},\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{B})^{\beta_{4}}+\beta_{5}\}\}$ acts as the constrained ratio of the trading fee to that net payment. We omit the parameters $\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5}$ in the rest of this paper for brevity.
5.3 Parameter Selection
@@
{restatable}
propostheoremnecessary Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, then the following conditions are necessary for $c$-smaller-better to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@
From the above theorem, we require that $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$. Next, we identify additional conditions that are sufficient for both properties to hold.
{restatable}
thmtheoremsufficient Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, if $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$ and $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$, then following items are sufficient for the $c$-Non-Splitting and $c$-smaller-better properties to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@
$\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\frac{R^{B}}{R^{A}}\times O^{A}\times\max\left\{r_{\min},\beta_{1}\times\frac{O^{A}}{R^{A}}+r_{\max}\right\}\,\,.$
By setting $\beta_{4}=1$ and $\beta_{5}=r_{\max}$ in the sufficient condition of the $c$-Non-Splitting and $c$-smaller-better properties (Theorem 5.3), the sufficient condition becomes: {restatable}cortheoremspecialpara For any $\beta_{1}<-1$ and $c$ satisfying
@@
6 Game-theoretic Analysis
The model gives rise to a game (§6.1) played among traders and liquidity providers. Then, we find a specific property that AMM algorithms should exhibit to enhance throughput (§6.2).
6.1 Game Model
Our model gives rise to a sequential game with discrete steps $k=0,1,2,\cdots$. The game is parameterized by the number $n$ of pools and by trading fee functions tf of the AMMs. We denote it by $\Gamma_{n}(\textit{tf})$.
6.1.1 System State
In $\Gamma_{n}(\textit{tf})$, the state of each pool $\textit{pool}_{i}$ consists of the amount of deposited token A, the amount of deposited token B and the amount of share tokens, $R^{A}_{i},R^{B}_{i},R^{S}_{i}$ respectively. Recall that share tokens are not deposited in the pool but are held by liquidity providers and $R^{S}_{i}$ is the total amount of its related share tokens held by liquidity providers. We denote the state of all AMM contracts in step $k$ by ${\textbf{R}({k})=\left(\left(R_{1}^{A}({k}),R_{1}^{B}({k}),R_{1}^{S}({k})\right),\cdots\left(R_{n}^{A}({k}),R_{n}^{B}({k}),R_{n}^{S}({k})\right)\right)}$.
6.1.2 Liquidity Provider Actions
The liquidity provider decides the amount of tokens she deposits in each AMM pool. We denote the amount of token A and token B depositted in $\textit{pool}_{i}$ by $l^{A}_{i},l^{B}_{i}\geq 0$, respectively. Recall that the scheduler assign the liquidity provider $l^{A}$ token A and $l^{B}$ token B. The total amount of tokens deposited should not exceed the amount she holds:
@@ 6.1.2 Liquidity Provider Actions
$\mathcal{A}_{lp}(l^{A},l^{B})=\\ \left\{a_{lp}\left|\forall 1\leq i\leq n,l^{A}_{i}=\frac{p^{B}}{p^{A}}l^{B}_{i}\geq 0,\sum_{i=1}^{n}l^{A}_{i}\leq l^{A},\sum_{i=1}^{n}l^{B}_{i}\leq l^{B}\right.\right\}\,\,.$
@@ 6.1.3 Trader Actions
Recall that after the trade operation and the arbitrage, $R^{A}_{i}$ and $R^{B}_{i}$ remain unchanged (Section 4.2). Consequently, the state of the liquidity pools remains unchanged in the subsequent step:
@@
6.1.4 Utility and Strategies
For traders and liquidity providers, we first discuss their revenue and then define their strategies and utility, respectively. We assume that traders and liquidity providers consider the value of tokens the same as the prices of the external market, namely $p^{A}$ and $p^{B}$.
@@ 6.1.4 Utility and Strategies
Denote the revenue of a liquidity provider with her action $a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)$, the action of the BA trader in the next step $a^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{n}\right)$ and the system state R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$, by the function $U_{lp}(\textbf{R},a_{lp},a^{BA})$. In the next step, $\textit{pool}_{i}$ receives a trading fee of $\textit{tf}(R^{A}_{i}+l^{A}_{i},R^{B}_{i}+l^{B}_{i},b^{\textit{BA}}_{i})$. The liquidity provider receives a fraction of that fee proportional to her fraction of share tokens out of all shares in the pool. Therefore, the revenue function is:
@@
6.1.5 Solution Concept
The subgame perfect Nash equilibrium (SPNE) ensures that in sequential games, players cannot gain higher utility by changing strategies at any game step, supported by backward induction [41] where players optimize their utility based on previous actions.
When a trader takes an actions in a given step, her utility is influenced solely by her immediate strategy and the current state of AMMs, as outlined in Equation 6.1.4. Crucially, future actions do not bear on this calculation allowing the trader to directly optimize her utility function, thereby establishing dominant strategies.
@@ 6.1.5 Solution Concept
Note that $\tau^{BA}$ is a best response for the BA trader. The strategy of liquidity provider in equilibrium is just the optimal strategy when traders adopt their best response, namely $\forall\textbf{R},l^{A},l^{B}$, we have
@@
6.2 Desired Property
@@
The strategy that uniformly at random selects an AMM contract to acquire all her needed token A is the perfect parallelism strategy: {restatable}defiperfectparallelismstrategy The perfect parallelism strategy of the BA trader is $\hat{\tau}^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$, where
@@
{restatable}
proppropertyperfectparallelism The perfect parallelism strategy of the BA trader is a dominant strategy, namely
@@
7 SAMM Equilibrium
We analyze the behavior of players in the game with SAMM. We first prove the trader randomly selects a pool to trade when the states of pools are balanced (§7.1). Then, we show the system tends to the balanced state since liquidity providers invest their tokens in the smallest pools, reducing the difference in the volume of pools (§7.2). Full proofs are in appendix G.
7.1 Trader Strategy
Consider the case that the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. As discussed in Section 5.3, the SAMM gross amount satisfies the $c$-non-splitting property and $c$-smaller-better property for a certain $10<c<$ 1. We assume that the required amount of token A, $b^{\textit{BA}}$, is at most a fraction $c$ of the amount of deposited token A in all pools, i.e.,
$\forall 1\leq i\leq n,b^{\textit{BA}}\leq cR^{A}_{i}\,\,.$
7.1.1 Traders’ optimal action
The $c$-non-splitting property and $c$-smaller-better property give a trader the incentive to randomly select one of the smallest pools to trade all her required tokens. Recall that $a_{i}^{\text{BA}}(b^{\textit{BA}})$ is the action of acquiring all $b^{\textit{BA}}$ token A in $\textit{pool}_{i}$ (Equation 20). We define the set of actions that trade in one of the smallest pools: {restatable}defismallestpoolactionset The Smallest Pool Action Set is the set of actions that acquire all $b^{\textit{BA}}$ token A in one of the smallest pools under state R:

### v2

@@ IV-C CPMM Trading fee
V SAMM: Sharded AMM
We introduce SAMM, the first sharded AMM. We explain its rationale and derive two desired properties of the contract (§V-A). We present its trading fee function (§V-B) and find parameters to fullfill both properties (§V-C).
V-A SAMM Structure and Properties
@@
{prop}
[$c$-Non-Spliting] Let $m\geq 2$. Given a set of output amount by $\{O^{A}_{j}|1\leq j\leq m,O^{A}_{j}>0\}$, denote by $\tilde{O}^{A}$ the sum of the amounts in the set, $\tilde{O}^{A}=\sum_{j=1}^{m}O^{A}_{j}$. For the constant $10<c<$ 1 and the deposited amount of tokens $R^{A}$ and $R^{B}$, if $\frac{\tilde{O}^{A}}{R^{A}}\leq c$ , then the cost of trading $\tilde{O}^{A}$ token A is less than the sum of the cost of trading $O^{A}_{j}$ token A for $1\leq j\leq m$, i.e.,
@@
V-A2 c-smaller-better property
Our second goal is to maintain balanced volumes in all shards. This is crucial because the volume directly affects the slippage. When there are stark differences in pool sizes, with some being much smaller than others, the slippage in trading within these smaller pools is significantly greater than in larger ones. This discrepancy can result in transactions clustering in the larger pools rather than spread evenly, reducing parallelism. Moreover, an attacker could disrupt the system by adding or removing liquidity from certain pools, thereby harming overall system performance.
We address this by incentivizing liquidity providers to allocate their tokens to the shards with lower volumes. Intuitively, this ensures that smaller pools receive more frequent fees from traders when the volumes of pools are not balanced, which incentivizes liquidity providers to deposit tokens in these smaller pools. Therefore, the system would converge to the state where all pools have balanced volumes, and traders then randomly select pools for trading. As with the $c$-Non-Splitting property, large transactions suffer from high slippage, resulting in an advantage for larger pools. So here too, we refine this requirement to scenarios where the ratio of the traded amount and the deposited amount is below a threshold $c$. We call this the $c$-smaller-better property.
{prop}
[$c$-smaller-better] Given an output amount ${O^{A}>0}$, for any two pools with deposited token amounts $(R^{A}_{i},R^{B}_{i})$ and $(R^{A}_{j},R^{B}_{j})$, respectively, and $R^{A}_{i}<R^{A}_{j}$. For the constant $10<c<$ 1, if $\frac{O^{A}}{R^{A}_{j}}<\frac{O^{A}}{R^{A}_{i}}\leq c$ and $\frac{R^{A}_{i}}{R^{B}_{i}}=\frac{R^{A}_{j}}{R^{B}_{j}}$, then the cost of trading $O^{A}$ token A in the smaller pool is less than that in the larger pool, i.e.,
@@ V-A3 $c$ value
The CPMM cost function, $G_{\textit{CPMM}}(R^{A},R^{B},O^{A})$, does not satisfy either property (Appendix C).
V-B Trading Fee Design
To satisfy properties V-A1 and V-A2, we first generalize the trading fee function to provide flexibility in the incentive design (§V-B1). Then, we propose a specific trading fee function (§V-B2).
V-B1 Generic Trading Fee Function
@@
V-B2 Bounded-Ratio Trading Fee Function
@@
$\textit{tf}(R^{A},R^{B},O^{A};\beta_{1},\beta_{2},\beta_{3},\beta_{4})\coloneqq\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}\,\,.$
@@
$\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A};\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5})\coloneqq\frac{R^{B}}{R^{A}}O^{A}\times\\ \max\{r_{\min},\min\{r_{\max},\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{A})^{\beta_{4}}+\beta_{5}\}\}$ (10)
The ratio $\frac{R^{A}}{R^{B}}$ represents the market price of token A relative to token B (see Section IV-B). The product $\frac{R^{B}}{R^{A}}O^{A}$ is thus the trader’s net payment in terms of token B without slippage. Then $\max\{r_{\min},\min\{r_{\max},\beta_{1}(R^{A})^{\beta_{2}}(R^{B})^{\beta_{3}}(O^{B})^{\beta_{4}}+\beta_{5}\}\}$ acts as the constrained ratio of the trading fee to that net payment. We omit the parameters $\beta_{1}$$\beta_{5}$ in the rest of this paper for brevity.
V-C Parameter Selection
@@
{propos}
[] Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, then the following conditions are necessary for $c$-smaller-better to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@
Based on Proposition V-C, we require that $\beta_{1}<0$, $\beta_{2}+\beta_{4}=0$, $\beta_{3}=0$, $0<\beta_{4}\leq 1$, and $r_{\min}<\beta_{5}\leq r_{\max}$. Next, we identify additional conditions that are sufficient for both properties to hold.
{thm}
[] Let $\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\textit{tf}_{\textit{BRP}}(R^{A},R^{B},O^{A})$, if $\beta_{1}<0,\beta_{2}+\beta_{4}=0,\beta_{3}=0,0<\beta_{4}\leq 1,r_{\min}<\beta_{5}\leq r_{\max}$ and $\frac{\beta_{5}-r_{\min}}{-\beta_{1}}\geq c^{\beta_{4}}$, then following items are sufficient for the $c$-Non-Splitting and $c$-smaller-better properties to hold for $G_{\textit{SAMM}}(R^{A},R^{B},O^{A})$:
@@
$\textit{tf}_{\textit{SAMM}}(R^{A},R^{B},O^{A})=\\ \frac{R^{B}}{R^{A}}\times O^{A}\times\max\left\{r_{\min},\beta_{1}\times\frac{O^{A}}{R^{A}}+r_{\max}\right\}\,\,.$
By setting $\beta_{4}=1$ and $\beta_{5}=r_{\max}$ in the sufficient condition of the $c$-Non-Splitting and $c$-smaller-better properties (Theorem V-C), the sufficient condition becomes: {cor}[] For any $\beta_{1}<-1$ and $c$ satisfying
@@
VI Game-theoretic Analysis
The model gives rise to a game (§VI-A) played among traders and liquidity providers. Our goal (§VI-B) is to ensure that at equilibrium, traders randomly select pools to trade in without splits, thereby enhancing throughput. SAMM achieves this goal (§VI-C) and converges to a state where all pools have balanced volumes, effectively resisting any attempts by attackers to unbalance the pool state.
VI-A Game Model
The model gives rise to a sequential game with discrete steps $k=0,1,2,\cdots$. The game is parameterized by the number $n$ of pools and by their trading fee function tf. We denote the game by $\Gamma_{n}(\textit{tf})$.
VI-A1 System State
In $\Gamma_{n}(\textit{tf})$, the state of each pool $\textit{pool}_{i}$ in step $k$ consists of the amount of deposited token A, the amount of deposited token B and the amount of share tokens, $R_{i}^{A}({k}),R_{i}^{B}({k}),R_{i}^{S}({k})$ respectively. Recall that share tokens are not deposited in the pool but are held by liquidity providers, so $R^{S}_{i}$ is the total amount of $\textit{pool}_{i}$’s share tokens held by liquidity providers. We denote by $\textbf{R}({k})$ the state of all AMM contracts in step $k$, ${\left(\left(R_{1}^{A}({k}),R_{1}^{B}({k}),R_{1}^{S}({k})\right),\cdots,\left(R_{n}^{A}({k}),R_{n}^{B}({k}),R_{n}^{S}({k})\right)\right)}$.
VI-A2 Liquidity Provider Actions
The liquidity provider decides the amount of tokens she deposits in each pool. Denote the amount of token A and token B she deposits in $\textit{pool}_{i}$ by $l^{A}_{i},l^{B}_{i}\geq 0$, respectively. Recall that the scheduler assign the liquidity provider $l^{A}$ token A and $l^{B}$ token B. The total amount of tokens deposited should not exceed the amount she holds:
@@ VI-A2 Liquidity Provider Actions
$\mathcal{A}_{lp}(l^{A},l^{B})=\left\{a_{lp}\left|\begin{aligned} &\forall 1\leq i\leq n,l^{A}_{i}=\frac{p^{B}}{p^{A}}l^{B}_{i}\geq 0,\\ &\sum_{i=1}^{n}l^{A}_{i}\leq l^{A},\sum_{i=1}^{n}l^{B}_{i}\leq l^{B}\end{aligned}\right.\right\}\,\,.$
@@ VI-A3 Trader Actions
Recall that after the trade operation and the arbitrage, $R^{A}_{i}$ and $R^{B}_{i}$ remain unchanged (Section IV-B). Consequently, the state of the liquidity pools remains unchanged in the subsequent step:
@@
VI-A4 Utility and Strategies
For traders and liquidity providers, we first discuss their revenue and then define their strategies and utility, respectively. Players determine the value of tokens according to the external market, namely $p^{A}$ and $p^{B}$.
@@ VI-A4 Utility and Strategies
Denote the revenue of a liquidity provider with her action $a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)$, the action of the BA trader in the next step $a^{\textit{BA}}=\left(b^{\textit{BA}}_{1},\cdots,b^{\textit{BA}}_{n}\right)$ and the system state R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$, by the function $U_{lp}(\textbf{R},a_{lp},a^{BA})$. In the next step, $\textit{pool}_{i}$ receives a trading fee of $\textit{tf}(R^{A}_{i}+l^{A}_{i},R^{B}_{i}+l^{B}_{i},b^{\textit{BA}}_{i})$. The liquidity provider receives a fraction of that fee proportional to her fraction of share tokens out of all shares in the pool. Therefore, the revenue function is
@@
VI-A5 Solution Concept
In a subgame perfect Nash equilibrium (SPNE), players cannot gain higher utility by changing strategies at any step [45], knowing subsequent players will take their best responses.
When a trader takes an action in a given step, her utility is influenced solely by her immediate strategy and the current state of AMMs, as outlined in Equation VI-A4. Crucially, future actions do not affect this calculation allowing the trader to directly optimize her utility, thereby establishing dominant strategies.
@@ VI-A5 Solution Concept
Note that $\tau^{BA}$ is a best response for the BA trader. The strategy of liquidity provider in equilibrium is just the optimal strategy when traders adopt their best response, namely for all R, $l^{A}$, and $l^{B}$, we have
@@
VI-A6 Game Assumptions
We briefly discuss two non-trivial model assumptions.
Gas fees
In a blockchain, each transaction consumes resources measured in a metric called gas. The transaction pays a fee to miners according to its gas consumption. In practice, gas fees are volatile, often nominal in high-performance blockchains (e.g., in Sui under \$$\$0.003$ per transaction on average). The model conservatively neglects gas fees to avoid arbitrary assumptions on gas costs. Gas fees disincentivize trade splitting, which helps with parallelization. In particular, with a set gas fee, Properties V-A1 and V-A2 continue to hold. Therefore, it would not change the equilibrium analysis.
Arbitrageurs
Perfect arbitrageurs are a common assumption in the literature [36, 24, 12, 13]. We nonetheless demonstrate later (Section VIII) that our results hold with real workloads, without this assumption.
VI-B Desired Property
@@
The strategy that uniformly at random selects an AMM contract to acquire all her needed token A is the perfect parallelism strategy: {defi}[] The perfect parallelism strategy of the BA trader is $\hat{\tau}^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$, where
@@
{prop}
[] The perfect parallelism strategy of the BA trader is a dominant strategy, namely
@@
VI-C SAMM Equilibrium
We analyze the SPNE of the above game. We only provide a roadmap here, from trader strategy to liquidity provider strategy, and defer the full analysis to Appendix G.
VI-C1 Trader Strategy
Consider the case that the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. As discussed in Section V-C, the SAMM gross amount satisfies the $c$-non-splitting property and $c$-smaller-better property for a certain $10<c<$ 1. We assume that the required amount of token A, $b^{\textit{BA}}$, is at most a fraction $c$ of the amount of deposited token A in all pools, i.e., $\forall 1\leq i\leq n,b^{\textit{BA}}\leq cR^{A}_{i}\,\,.$
The $c$-non-splitting property and $c$-smaller-better property give a trader the incentive to randomly select one of the smallest pools to trade all her required tokens. Recall that $a_{i}^{\text{BA}}(b^{\textit{BA}})$ is the action of acquiring all $b^{\textit{BA}}$ token A in $\textit{pool}_{i}$ (Equation 20). We define the set of actions that trade in one of the smallest pools: {defi}[] The Smallest Pool Action Set is the set of actions that acquire all $b^{\textit{BA}}$ token A in one of the smallest pools under state R: